# WikiGap: Promoting Epistemic Equity by Surfacing Knowledge Gaps Between English Wikipedia and other Language Editions

### v3

@@ 5. Findings
Despite minor confusion about the color scheme, most users described the overall experience as more enjoyable and cognitively less demanding than their usual experience with Wikipedia’s default ILL. P9 summarized this sentiment: “Sometimes even when I open my Wikipedia, I would not read that much because of laziness, but WikiGap did a good job extracting information and makes the whole experience much more pleasant.”
WikiGap Enhanced Fact-Finding Performance and Cross-Cultural Learning via Passive Discovery and Active Exploration (D4, D5)
@@
Figure 7. Average completion time (left) and quiz accuracy (right) across conditions. Blue represent the control condition (no WikiGap), and orange represent the treatment condition (with WikiGap). Using WikiGap demonstrates a statistically significant improvement in user performance on the quiz.
@@
Beyond performance gains, the extension also facilitated passive discovery and cross-cultural learning – a central goal of our design. Several users remarked that they encountered information they would not have actively sought out on their own. As P1 reflected, “Without [WikiGap], I wouldn’t have thought about looking into Wikipedia pages in different languages.” Others described moments of unexpected cultural insight, such as P5, who noted, “I would have completely missed out on the historical aspect if I used Wikipedia normally (only in English).” Participants also pointed to surprising culture-specific facts, like the culinary use of Sprite in a traditional dish (P13). These examples illustrate that WikiGap not only supported successful task completion but also broadened users’ awareness of culturally situated knowledge that is often absent from English Wikipedia.
WikiGap Raises Awareness of Multilingual Gaps and Motivates Exploration
@@
This realization often triggered a sense of frustration, as users became aware of the limits of English Wikipedia.
“I’m frustrated that this doesn’t already exist! I feel like I’ve been missing out on information from other languages because I just assumed other pages would have the same information translated.” — P5
Other participants had deeper reflections on how WikiGap changed their perception about knowledge: “Knowledge is not plain, knowledge has different dimensions waiting to be discovered, and this tool offers this opportune moment.” (P10).
Importantly, the awareness raised by WikiGap also led to changes in participants’ attitudes and future behaviors. Now they can “see different viewpoints from different language pages“ (P17), they are more inclined to explore content outside their primary language in the future. Several users described feeling more motivated to engage with multilingual content now that they understood its value.
@@
Echoing this, Since WikiGap displays multilingual content in one centralized place, it provides great convenience to explore cross-lingual knowledge, and contributes to users’ willingness to engage. As P1 noted, “If there’s a tool like WikiGap that can help me easily get various information from other pages, I would like to explore. If not, I wouldn’t navigate multiple pages myself.”
④Respondents’ Views on LLMs and Wikipedia: Perceived Trust and Tensions
@@
WikiGap Supports Editors in Validating and Integrating Multilingual Content
@@
Some users noted that the ability to trace and verify information from other language editions made the tool useful for improving article quality. As one frequent editor explained, “Sometimes an article in another language has crucial information that’s missing in English. WikiGap helps me find and translate it, then cite it properly.” (P16). Another participant elaborated on how the tool fits into their existing editing workflow: “I’d check the source WikiGap shows me, then cite the same source… WikiGap helps me discover facts and references.” (P17). While these reflections were not the result of targeted prompts or interview questions targeted towards editing-specific tasks, they highlight WikiGap’s broader potential to support contributors in bridging cross-lingual knowledge gaps.
@@ 6. Discussion
In this work, we sought to combat the epistemic injustice (Ajmani et al., 2023a) posed by the dominance of the English Wikipedia project over other language editions. We did so by leveraging novel affordances brought by advances in fine-grained cross-linguistic comparison (Samir et al., 2024), enabling us to identify factual gaps in English Wikipedia and succinctly display them to readers. By surfacing complementary facts through WikiGap, our system performs an infrastructural inversion (Star, 1999): it makes visible the hidden cross-lingual knowledge gaps while productively challenging the long-standing English-as-superset assumption. Importantly, this intervention is introduced with minimal disruption to the familiar English Wikipedia reading experience. Beyond enabling access to a more complete set of factual information, our work educates readers about how knowledge on multilingual Wikipedia is composed and contested. Through this sociotechnical interaction, WikiGap demonstrates how systems can foster more equitable collaboration in distributed knowledge production across linguistic boundaries (Ackerman, 2000), reshaping how knowledge is consumed and shared across communities. Below, we discuss how English Wikipedia came to have a limited demographic representation in its editor base, yet an outsized global readership (Sec. 6.1); how the English-as-superset hypothesis still persists among readers, despite its debunking in prior seminal studies from over a decade ago (e.g., Hecht and Gergle, 2010) (Sec. 6.2); and why this epistemology is a byproduct of the subsuming culture of computationalism (Golumbia, 2009), which (in part) explains its stubborn persistence (Sec. 6.3).
6.1. Editor-Readership Asymmetry in English Wikipedia
While English has undeniably achieved the status of a global lingua franca (Schneider, 2019), this does not mean that English speakers worldwide have had equal opportunity to contribute to English Wikipedia. Wikipedia as a whole is animated by an ethos of participatory, democratic knowledge curation. Yet from its inception, historical inequities have undermined this libertarian ideal (Shaw and Hargittai, 2018). Participation in the internet has been unequal, both within and across nation-states, with greater access among socially privileged individuals. Four years after English Wikipedia was launched, access to the internet for developing countries was still in the single digits, percentage-wise (Chen and Wellman, 2005). Within nation states, internet usage is clearly stratified across gender, socioeconomic class, race, and a stark urban-rural divide (Chen and Wellman, 2005), many of which have also been studied in the context of English Wikipedia (Johnson et al., 2016; Hargittai and Shaw, 2015; Adams et al., 2019). The inequities that were present during the establishment of English Wikipedia matter, as peer production processes are understood to ossify and concentrate power among a small group of early editors, resembling an oligarchy (Shaw and Hill, 2014). This editor–reader asymmetry exemplifies a socio-technical gap (Ackerman, 2000), in which the ideals of open participation clash with the realities of unequal collaboration in a large-scale knowledge infrastructure.
Although these disparities are well studied, Wikimedia has struggled to reduce these editor demographic skews. As of (2024b), English Wikipedia contributions are still overwhelmingly contributed by male-identifying editors, and there is a remarkable lack of Black and Hispanic editors from within the United States (Wikimedia Foundation, 2024b).
Wikipedia’s most prominent pillar, the Neutral Point-of-View (NPOV) policy, has been roundly criticized for obstructing more equitable production of knowledge. In particular, it runs counter to feminist epistemology that the act of knowing is necessarily situated, that there is no voice from nowhere (Haraway, 2013). Importantly, the policy has been criticized for obscuring the role of patriarchic and Anglocentric power structures in how knowledge produced by different communities is appraised (Foucault, 2013; Kumar, 2017). Kumar (2017) situates this discourse in a case-study on Indian editors presenting arguments on the naming of the Ganga river (as opposed to the Ganges river) on English Wikipedia. Ultimately, the smaller group of Indian editors was overwhelmed by the preponderance of reliable Anglocentric sources that asserted the veracity of the Ganges label. The case-study is illustrative of how the notion of neutrality is in fact biased towards information that is disseminated from Anglocentric institutions. Importantly, this discounts against knowledge derived from lived experiences and societies that privilege oral traditions, or have had historically unequal opportunities in access to the internet (Avieson and Mazumdar, 2025).
It is important to note that English Wikipedia attracts a massive readership, even outside of the geographic Anglosophere. As Kumar (2017) notes (at the time), $20\%$ of English Wikipedia’s visitors were based in India, even more than readers from the United Kingdom. The composition of English Wikipedia’s editor base is thus hardly representative of the broader population of literate, internet-equipped English speakers. As acknowledged in responses to our open-ended surveys, readers believe in the knowledge of local authorities. As one participant noted: “diverse cultural perspectives on food is more linked to the cultural ecosystem where the food is grounded” (P7). Indeed, the capacity for volunteers to contribute and disseminate information from geographically-situated sources was seen as one of the primary benefits of Wikipedia (Sen et al., 2015), Yet prior work highlights that readers are not simply passive consumers but active participants with distinct roles and needs, and that their preferences do not always align with where editorial effort is concentrated. This production–consumption gap poses a fundamental challenge for distributed collaborative work: designing socio-technical processes of knowledge creation that account for readers as well as editors (Antin and Cheshire, 2010; Lehmann et al., 2014). There is thus a strong impetus to achieve more equitable knowledge representation. WikiGap presents an important contribution towards achieving greater epistemic equity.
6.2. English-as-superset Hypothesis Lives On
Although there are well-documented structural barriers to contributing to English Wikipedia, its overwhelming dominance makes it appear as a superset of the information contained in all other language editions, rather than a necessarily biased and socio-geographically-positioned construction (Kumar, 2021). Hecht (2013) studied this perception of English Wikipedia, calling it the English-as-superset hypothesis, writing that the hypothesis is “commonly adopted in Wikipedia-based research projects and systems”. Practically, this has entailed designating English Wikipedia as the ground-truth language edition for assessing, for example, the correctness of question-answering systems (Asai et al., 2021).
Although there is now greater appreciation among Wikipedia researchers of the cross-linguistic information diversity between Wikipedia language editions, there is little reason to expect the broader Wikipedia userbase outside of Wikipedia experts to be privy to these results. As Hickman et al. (2021) documented in their cross-linguistic comparative analysis of articles pertaining to Kashmir, editors of Hindi and Urdu articles drew on English Wikipedia for their primary research to a large extent. Their editing process thus implies an underlying acceptance of the English-as-superset hypothesis, where English Wikipedia is perceived as neutral, objective, and, critically, complete.
Our user studies also demonstrate the prevalence of the hypothesis, with some participants commenting on the hypothesis directly: “Before WikiGap, I always assumed that everything presented in English would also be visible in another language.” Thus, while Hecht and Gergle (2010) effectively critiqued the English-as-superset hypothesis $15$ years ago, its continued prevalence calls for greater institutional intervention to increase awareness among the broader Wikipedia userbase of the vast cross-linguistic diversity between language editions.
Our research demonstrates that the WikiGap interface directly challenges the English-as-superset perception as well as the invisibility problem: The sentence-level complementary facts presented by WikiGap function as boundary objects (Star and Griesemer, 1989), serving users across different linguistic communities by making the otherwise hidden contributions of diverse editor communities visible and facilitating understanding of the distributed nature of collaborative knowledge production. As such, adoption of our interface can challenge the problematic status quo.
6.3. Objectivity in Knowledge Representation
It should be acknowledged however that the epistemic assumption of the English-as-superset hypothesis is deeply ingrained in the broader culture of computationalism (Golumbia, 2009). Computationalism is concerned with treating the mind as a computer, and thereby entails the aim of codifying knowledge with digital precision. This entails that all concepts, even socioculturally-contingent artifacts like cultural dishes, have one-to-one definitions, in the form of a single collection of facts (Golumbia, 2009, Chapter 6). The role of language here, then, is a cryptographic encoding scheme, rather than a reflection of socio-geographic context (Golumbia, 2009, Chapter 4). Within this cultural philosophy, it makes sense that cross-linguistic diversity between language editions are seen by some as a deviation, a shortcoming, relative to a monolithic, consistent set of facts. This is why some computational research projects have formulated cross-linguistic diversity as a problem to be solved (see Sec. 2).
Our study showed that some participants, although a minority, held perspectives that are aligned with computationalism (Sec. 5). That is, they viewed the gaps in English Wikipedia (foregrounded by WikiGap) as a shortcoming of the English language edition, making them perceive it as less “reliable” and “complete”, and even asking, akin to the work we describe in Sec. 2, why all Wikipedia language editions don’t contain the same set of facts. We thus see that cross-linguistic diversity is perceived by some as being in opposition to neutrality and objectivity. One of these participants even said they would prefer to turn to LLMs, which conveys what they perceive as a consolidated view of all information on a prompted topic. In this light, we can understand that some view LLMs as the purest form of the “voice from nowhere” (Haraway, 2013), in practically a literal sense.
While there have been anxieties circulating around how contributions to Wikipedia might be affected by the emergence of LLMs (Maiberg, 2024), our work highlights that Wikipedia’s role in presenting information to readers is similarly in flux. That is, our work demonstrated that readers looking for an (ostensibly) objective summary of a topic may turn to LLMs over Wikipedia. While LLMs pose some advantages (directly answering user queries), they also marginalize the role of provenance in accessing information (Shah and Bender, 2022). WikiGap presents one approach for encouraging continued readership of Wikipedia as it highlights one of its foundational strengths: an established platform for soliciting citation-backed Volunteered Geographic Information from all around the world (Sen et al., 2015), containing a plurality of socio-geographically positioned perspectives (Figure 3). The current multilingual reading experience facilitated by Interlanguage Links fails to foreground this strength, however. Instead, readers are led to believe that Wikipedia’s language editions reflect global consensus – that the information on the pages are consistent. As Hecht and Gergle (2010) explain, a natural corollary of believing in global consensus is the English-as-superset hypothesis, which contributes a form of epistemic injustice by marginalizing novel information in other language editions. As demonstrated in our user evaluations, WikiGap does well to promote epistemic equity across language editions.
@@ 6.4. Limitations and Future Work
In this study, we intentionally focused on English and three additional language editions, namely Russian, French, and Chinese. We chose the first two as the performance of InfoGap was validated on English, French, and Russian (Samir et al., 2024), and we additionally evaluated the performance of the system on Chinese for this work as some co-authors were proficient in Chinese. Several participants expressed a desire for broader language coverage, particularly for pages in their language of origin. For example, one participant noted, “I would love to see how the Tagalog page could complement it. In fact, diverse cultural perspectives on food are more linked to the cultural ecosystem where the food is grounded” (P7). Expanding to more languages will enrich cross-cultural perspectives and align with users’ expectations in future work. The underlying InfoGap pipeline is language-agnostic: in principle, it can operate on any language that GPT-4 supports, though quality may vary, especially for low-resource languages. This is important for future research as smaller language editions remain largely understudied and present additional challenges in their own right (Nigatu et al., 2024; Khatri et al., 2022).
@@
Other barriers in facilitating epistemic justice in Wikipedia.
WikiGap presents only one potential component of the puzzle towards achieving greater epistemic justice in Wikipedia. But other critical challenges have been identified in facilitating the growth of smaller language editions. To name a few, the gender divide that has been reported largely in English Wikipedia is present in many other language editions (Johnson et al., 2021). Some smaller language editions are difficult to contribute to due to the lack of language technology support that is taken for granted in larger ones, such as spellcheck and translation (Nigatu et al., 2024). Other language editions lack scholarly sources in the local language, making it difficult to corroborate contributed information (Nigatu et al., 2024; Sen et al., 2015). Smaller language editions also lack established mentorship structures, making it difficult for new editors to contribute content (Khatri et al., 2022). All of these challenges ultimately affect WikiGap, as they limit the set of useful facts that can be presented in the interface.

### v4

@@ 1. Making Multilingual Differences Visible Improves User Preference and Reading Experience (D1, D2, D3)
Despite minor confusion about the color scheme, most users described the overall experience as more enjoyable and cognitively less demanding than their usual experience with Wikipedia’s default ILL interface. P9 summarized this sentiment: “Sometimes even when I open my Wikipedia, I would not read that much because of laziness, but WikiGap did a good job extracting information and makes the whole experience much more pleasant.”
2. Visibility of Cross-Lingual Differences Improves Fact-Finding Performance and Supports Cross-Cultural Learning (D4, D5)
@@
Figure 7. Average completion time (left) and quiz accuracy (right) across conditions. Blue represents the control condition (no WikiGap), and orange represents the treatment condition (with WikiGap). Using WikiGap demonstrates a statistically significant improvement in user performance on the quiz.
@@
Beyond performance gains, the extension also facilitated passive discovery and cross-cultural learning – a central goal of our design. Several users remarked that they encountered information they would not have actively sought out on their own. As P1 reflected, “Without [WikiGap], I wouldn’t have thought about looking into Wikipedia pages in different languages.” Others described moments of unexpected cultural insight, such as P5, who noted, “I would have completely missed out on the historical aspect if I used Wikipedia normally (only in English).” Participants also pointed to surprising culture-specific facts, like the culinary use of Sprite in a traditional dish (P13). These examples illustrate that making multilingual differences visible not only improved task efficiency, but also broadened users’ exposure to culturally situated knowledge.
3. WikiGap Raises Awareness of Multilingual Gaps and Motivates Exploration
@@
This realization often prompted frustration as participants became aware of the limits of English Wikipedia. As P5 expressed, “I’m frustrated that this doesn’t already exist! I feel like I’ve been missing out on information from other languages because I just assumed other pages would have the same information translated.” Other participants had deeper reflections on how WikiGap changed their perception about knowledge: “Knowledge is not plain, knowledge has different dimensions waiting to be discovered, and this tool offers this opportune moment.” (P10).
Importantly, increased awareness also motivated future engagement. Participants described being more inclined to explore content outside their primary language now that differences were visible and easy to access.
@@
Several emphasized that centralizing multilingual content within a single interface reduced the effort required to engage, making exploration more likely. Now they can “see different viewpoints from different language pages“ (P17), they are more inclined to explore content outside their primary language in the future. Several users described feeling more motivated to engage with multilingual content now that they understood its value. As P1 noted, “If there’s a tool like WikiGap that can help me easily get various information from other pages, I would like to explore. If not, I wouldn’t navigate multiple pages myself.”
4. Respondents’ Views on LLMs and Wikipedia: Perceived Trust and Tensions
@@
5. WikiGap Suggests Opportunities for Supporting Cross-Lingual Editing
@@
Some users noted that the ability to trace and verify information from other language editions made the tool useful for improving article quality. As one frequent editor explained, “Sometimes an article in another language has crucial information that’s missing in English. WikiGap helps me find and translate it, then cite it properly.” (P16). Another participant elaborated on how the tool fits into their existing editing workflow: “I’d check the source WikiGap shows me, then cite the same source… WikiGap helps me discover facts and references.” (P17). While these reflections were not prompted by editor-specific questions, they point to an emergent opportunity: making multilingual gaps visible to readers may also lower barriers to cross-lingual knowledge integration, extending WikiGap’s impact beyond consumption to potential curation.
@@ 6. Discussion
We interpret our findings through the lens of boundary objects—shared artifacts that enable coordination across heterogeneous communities without requiring consensus or uniform interpretation (Star and Griesemer, 1989; Star, 2010). We argue that WikiGap functions as a boundary object within Wikipedia’s multilingual ecosystem by making cross-lingual knowledge differences legible, actionable, and traceable within everyday reading practices. In this role, WikiGap enables three intertwined forms of epistemic and coordinative work. First, it reframes readers’ epistemic assumptions by exposing the incompleteness of English Wikipedia and foregrounding knowledge produced by other language communities (§6.1). Second, it provides a shared reference point that supports articulation work between readers and editors across language editions, enabling verification and coordination without enforcing uniformity (§6.2). Third, it gestures toward a broader infrastructural role by modeling how pluralistic, provenance-aware interfaces can counter knowledge consolidation in emerging AI-mediated systems (§6.3).
6.1. Promoting Epistemic Equity by Challenging the “English-as-Superset” Assumption
Our user study revisits prior findings that the dominance of English Wikipedia leads readers to treat it as a comprehensive superset of knowledge across language editions, rather than as a culturally situated and socio-geographically positioned construction (Kumar, 2021). Participants explicitly articulated this belief. Although more than 60% of participants reported a non-English primary language, all identified English as their primary source on Wikipedia, largely because English Wikipedia covers more topics and is updated more frequently. Despite being familiar with Wikipedia’s ILLs, participants predominantly consumed content in English, reflecting how infrastructural dominance concentrates reader attention within a single language edition and reinforces the invisibility of knowledge produced by other language communities. At the interface level, Wikipedia’s current design renders other language editions technically accessible but epistemically invisible: ILLs signal the existence of alternative versions, yet do not communicate that English content is neither exhaustive nor neutral. This gap was evident in participants’ assumptions; as one remarked, “Before WikiGap, I always assumed that everything presented in English would also be visible in another language.”
Consistent with this perception, participants strongly preferred WikiGap over the default ILLs interface and performed significantly better on fact-finding tasks, indicating that access alone—without cues about epistemic differences, provenance or community context—is insufficient to support meaningful cross-lingual engagement. WikiGap challenges the “English-as-superset” assumption by both raising awareness of cross-linguistic knowledge gaps and redistributing reader attention toward knowledge produced by other language communities. Adopting an asset-based design orientation, the system foregrounds non-English content not as supplementary or corrective, but as a valuable product of distinct peer production efforts.
By embedding sentence-level complementary facts in context and pairing them with explicit provenance, WikiGap makes the limits of English Wikipedia visible and simultaneously inviting engagement with the work of other language editions. This design encourages readers to recognize Wikipedia as a collection of interconnected yet distinct knowledge production communities, rather than a single monolithic source of global consensus. In doing so, WikiGap promotes epistemic equity in a scoped, design-operational sense: by redistributing attention and engagement across Wikipedia’s multilingual peer production communities and supporting cross-community interaction as part of everyday readership.
6.2. WikiGap as a Boundary Object Supporting Articulation Work Across Reader and Editor Communities
Wikipedia’s multilingual knowledge production spans heterogeneous communities with distinct linguistic, cultural, and editorial norms. Our findings show that WikiGap operates as a boundary object in practice, providing a shared point of reference through which readers and editors from different language communities can interpret, evaluate, and act on cross-lingual knowledge differences without requiring consensus. By surfacing sentence-level complementary facts with clear provenance, WikiGap offers a way to support shared understanding across communities without requiring consensus or uniformity. Building on these findings, we interpret WikiGap as enabling an asset-based stance: it helps readers and multilingual contributors recognize non-English facts as usable resources by making them legible and actionable through provenance and in-context integration.
6.2.1. For Readers of English Wikipedia.
WikiGap benefits readers by supporting fact-finding, cross-cultural learning, and interpretive flexibility. Participants described discovering content they “would have otherwise missed” in English and expressed surprise at how much meaningful information existed exclusively in other language editions (Findings 1–2). Several noted that the tool increased their curiosity and willingness to visit other language pages, shifting their sense of Wikipedia from a monolithic source of global consensus to a collection of situated, culturally inflected knowledge communities.
Through a CSCW lens, this suggests that WikiGap enhances the legibility of multilingual knowledge work. It exposes facts that English readers typically never encounter, foregrounds the situated perspectives of marginalized linguistic communities, and supports interpretive flexibility—one of the core affordances of boundary objects. Rather than collapsing perspectives into a single authoritative narrative, WikiGap allows readers to compare, contextualize, and reinterpret information based on their own background knowledge and cultural standpoint.
6.2.2. For (Potential) Editors Across Languages.
Although our study focused on readers, participants with editing experience highlighted how WikiGap could support multilingual editorial workflows, they pointed to using the provenance link to see the original facts from the other languages to verify and potentially transfer content across editions. Maintaining consistency across language editions requires editors to reconcile differing source traditions, cultural emphases, and update rhythms. This process is inherently distributed and involves articulation work (Schmidt and Bannon, 1992): interpreting discrepancies, assessing their relevance to local communities, coordinating with other editors, and making decisions about whether and how to incorporate missing knowledge.
Participants noted that WikiGap reduces some of this articulation burden by automatically surfacing contextualized discrepancies, highlighting their provenance, and linking directly to original sentences. These cues help editors evaluate the origins and significance of differences, understand when they reflect meaningful cultural divergence versus simple incompleteness, and coordinate potential updates. Rather than automating agreement or enforcing uniformity, WikiGap supports the human work of interpretation and negotiation that underlies Wikipedia’s collaborative ecosystem. By enabling multiple situated viewpoints to coexist while still facilitating coordination across them, the system exemplifies how boundary objects can scaffold the complex articulation work required for distributed knowledge production.
In this way, WikiGap renders the often invisible coordinative labor of multilingual maintenance more visible and manageable by making cross-lingual differences legible and actionable through provenance and in-context integration, reducing coordination friction and supporting editor engagement.
6.3. Positionality, LLMs, and the Future Role of WikiGap as Boundary Infrastructure
Across both quantitative preferences and qualitative interviews, participants expressed strong appreciation for WikiGap’s explicit display of provenance and its framing of multilingual knowledge as situated rather than objective. This recognition of positionality became particularly salient when contrasted—often spontaneously—against the increasingly common practice of using Large Language Models (LLMs) as ostensibly objective, consolidated knowledge sources.
A small subset of participants (n=2) expressed a desire for consolidated knowledge and therefore perceived gaps highlighted by WikiGap as evidence of English Wikipedia’s “incompleteness.” One participant noted that they would prefer to turn to LLMs, which present a single synthesized output. While LLM outputs are often perceived as authoritative or persuasive (Chockkalingam et al., 2025), they rarely foreground multiple cultural or linguistic perspectives and frequently exhibit English-centric biases (Hershcovich et al., 2022; Mihalcea et al., 2025; Tao et al., 2024). They also obscure provenance (Shah and Bender, 2022), making it difficult for users to evaluate where information originates. LLMs thus risk presenting a “view from nowhere,” (Haraway, 2013) even though their outputs reflect specific training data and design choices.
Although WikiGap and LLMs appear to sit at opposite ends of the spectrum—positionality versus objectivity—they need not be mutually exclusive. Our findings suggest that asset-based design principles foregrounding positionality can inform future LLM interfaces by encouraging users to see responses as situated rather than uniformly objective. Concretely, interfaces could surface multiple culturally grounded viewpoints instead of a single synthesized answer, along with lightweight provenance cues indicating which language communities or traditions each perspective reflects. Such designs would counter the tendency of current systems to amplify majority perspectives and obscure minority ones, helping users explore diverse interpretations more intentionally. They would also make latent cultural knowledge in LLMs more accessible to researchers and end users alike. In this sense, embedding WikiGap-like scaffolding into LLM interfaces represents a promising direction for foregrounding epistemic positionality within increasingly consolidated AI ecosystems. This extends our system’s asset-based design principles to cross-lingual and multilingual LLM interfaces, encouraging users to engage with knowledge as situated, grounded in provenance, and produced by distinct language communities.
@@ 6.4. Limitations and Future Work
Our study focuses on English and three target languages (French, Russian, and Chinese). While these languages span different linguistic families and large Wikipedia communities, they do not capture the full diversity of Wikipedia’s language ecosystem. Several participants expressed a desire for broader language coverage, particularly for pages in their language of origin. For example, one participant noted, “I would love to see how the Tagalog page could complement it. In fact, diverse cultural perspectives on food are more linked to the cultural ecosystem where the food is grounded” (P7). Expanding to more languages will enrich cross-cultural perspectives and align with users’ expectations in future work. The underlying InfoGap pipeline is language-agnostic: in principle, it can operate on any language that GPT-4 supports, though quality may vary, especially for low-resource languages. This is important for future research as smaller language editions remain largely understudied and present additional challenges in their own right (Nigatu et al., 2024; Khatri et al., 2022).
@@
Future Work.
Building on our discussion of pluralistic LLM interfaces, future work could extend WikiGap’s design principles to other multilingual information-sharing platforms. Prior work shows that platforms such as Twitter and Flickr also exhibit substantial cross-lingual divergence in how events, entities, and cultural topics are discussed (Graham et al., 2014; Mendelsohn et al., 2023; Dotan and Zaphiris, 2010; Dong and Fu, 2010). Applying WikiGap could help users interpret how narratives differ across languages and communities beyond Wikipedia. Extending this approach across platforms would also inform the design of more culturally aware LLM interfaces, reinforcing our broader argument that pluralistic system design can mitigate the homogenizing effects of consolidated AI infrastructures.